# Parameter-Efficient Fine-Tuning in Spectral Domain for Point Cloud Learning

### v1

@@ Abstract
Recently, leveraging pre-training techniques to enhance point cloud models has become a hot research topic. However, existing approaches typically require full fine-tuning of pre-trained models to achieve satisfied performance on downstream tasks, accompanying storage-intensive and computationally demanding. To address this issue, we propose a novel Parameter-Efficient Fine-Tuning (PEFT) method for point cloud, called PointGST (Point cloud Graph Spectral Tuning). PointGST freezes the pre-trained model and introduces a lightweight, trainable Point Cloud Spectral Adapter (PCSA) to fine-tune parameters in the spectral domain. The core idea is built on two observations: 1) The inner tokens from frozen models might present confusion in the spatial domain; 2) Task-specific intrinsic information is important for transferring the general knowledge to the downstream task. Specifically, PointGST transfers the point tokens from the spatial domain to the spectral domain, effectively de-correlating confusion among tokens via using orthogonal components for separating. Moreover, the generated spectral basis involves intrinsic information about the downstream point clouds, enabling more targeted tuning. As a result, PointGST facilitates the efficient transfer of general knowledge to downstream tasks while significantly reducing training costs. Extensive experiments on challenging point cloud datasets across various tasks demonstrate that PointGST not only outperforms its fully fine-tuning counterpart but also significantly reduces trainable parameters, making it a promising solution for efficient point cloud learning. More importantly, it improves upon a solid baseline by +2.28%, 1.16%, and 2.78%, resulting in 99.48%, 97.76%, and 96.18% on the ScanObjNN OBJ_BG, OBJ_OBLY, and PB_T50_RS datasets, respectively. This advancement establishes a new state-of-the-art, using only 0.67% of the trainable parameters. The code will be made available at https://github.com/jerryfeng2003/PointGST.
@@ 1 Introduction
Point cloud learning is a fundamental and highly practical task within the computer vision community, drawing significant attention due to its extensive applications such as autonomous driving [31, 13], 3D reconstruction [45, 80], and embodied intelligence [82], among others. Different from images, point clouds are inherently sparse, unordered, and irregular, posing unique challenges to effective analysis.
@@
Nowadays, exploring pre-training techniques to boost the models has become a hot research topic in both natural language processing [9] and computer vision [21], and has been successfully transferred to the point cloud domain [36, 47, 85, 75]. For example, representative works like Point-BERT [85] and Point-MAE [47] introduce the Mask Point Modeling task for point cloud, and the subsequent studies [11, 93, 4, 36] propose effective customized pre-training tasks. After pre-training, these methods adopt the classical fully fine-training (FFT) strategy, where all models’ parameters are fine-tuned, leading to notable performance improvements and faster convergence compared to training from scratch. However, the FFT incurs significant computational costs, particularly in terms of GPU memory and storage, since fine-tuning the entire pre-trained models involves updating a large number of parameters. With the scaling of models [^1] Pre-trained point cloud model parameters increased 30$\times$ in three years, from 22.1M in ECCV 2022 [47] to 657.2M in ECCV 2024 [51]. or increasing needs to fine-tuning enormous new datasets, the demand for storing fine-tuned checkpoints grows, resulting in heightened storage and memory consumption.
@@
The community is well aware of this issue, and to mitigate this challenge, research efforts [88, 99, 64, 15] have been put on looking for a promising fine-tuning strategy, i.e., Parameter-Efficient Fine-Tuning (PEFT). Their common idea is to freeze the pre-trained models and simply insert extra learnable parameters in the models. The pioneers are IDPT [88] and DAPT [99], taking an important step toward efficient tuning on point cloud tasks. Specifically, IDPT [88] proposes an instance-aware dynamic prompt tuning to generate a universal prompt for the point cloud data. DAPT [99] generates a dynamic scale for each point token, considering the token significance to the downstream task.
@@
Figure 1: (a) The comparison of existing PEFT methods [99, 88] and our method. Instead of spatial tuning, we transform the point cloud to the spectral domain to mitigate the confusion among tokens and bring the intrinsic information of the downstream point clouds for targeted tuning. (b) The proposed PointGST outperforms all previous point cloud analysis methods while using extremely few trainable parameters. We even achieve a remarkable accuracy of 99.48% on the popular ScanObjectNN OBJ_BG dataset, pushing the performance on this dataset to surpass 99% for the first time.
@@
While the above methods have demonstrated efficiency in terms of training and storage, they still fail to achieve satisfactory performance across various pre-trained models. We attribute this to two-fold: First, essentially, these methods are fine-tuned in the spatial domain (Fig. 1(a)). In the PEFT setting, pre-trained tasks are designed to learn general representations of point clouds and lack downstream tasks prior. This leads to the inner tokens (features) from frozen pre-trained models that struggle to distinguish the fine-grained structures of the point cloud, where such tokens are referred to as inner confused tokens. Current point cloud PEFT methods fine-tune in the spatial domain by merging inner confused tokens with newly introduced learnable modules for downstream task adaptation, making it challenging to de-correlate the inherent confusion of the tokens. Second, intrinsic structures (e.g., task-specific geometry and relationship of points) of the downstream point clouds are essential for comprehensive analysis [80, 17]. The fixed pre-trained models lack the ability to update their parameters to learn the intrinsic information, relying solely on the representations captured during pre-training. Existing point cloud PEFT methods mainly utilize the encoded features from the frozen models to generate prompts or feed into adapters, which do not explicitly introduce the intrinsic information of the downstream point clouds.
@@
Considering the above issues, we shift our focus to a new perspective: spectral domain fine-tuning. Spectral representations offer significant distinctness over spatial domain representations. Specifically, compared to the spatial domain, features in the spectral domain can be easily de-correlated [57, 90], due to the ability of spectral representations to disentangle complex spatial relationships into distinct frequency components. This decorrelation mitigates the confusion among tokens by using orthogonal components (spectral basis) to separate them. Additionally, the spectral basis obtained by the original point cloud offers direct intrinsic information about the downstream tasks to better guide the fine-tuning process. However, effectively and efficiently fine-tuning the parameters in the spectral domain is not a trivial problem, where two crucial concerns need to be considered: 1) what spectral transform to use, and 2) how to inject the intrinsic information into the fixed pre-trained model by using this transform. An ideal transform for point clouds should be easy to construct, preserve intrinsic information, and adapt to the irregular structure of point cloud data. We find that the Graph Fourier Transform (GFT) stands out as particularly suitable. A properly constructed graph can provide a natural representation of the irregular point clouds that are adaptive to their structure, and the graph spectral domain can explicitly reveal geometric structures, from basic shapes to fine details, within a compressed space [25].
@@
In this paper, we explore the potential of fine-tuning parameters in the spectral domain. Our approach, illustrated in Fig. 1(a), introduces a novel PEFT method that optimizes a frozen pre-trained point cloud model through Graph Spectral Tuning, which we refer to as PointGST. Specifically, we first construct a series of multi-scale point cloud graphs that enable the calculation of global and local spectral basis, which will bring the intrinsic information like the geometry of the original point cloud. Then, we propose a lightweight Point Cloud Spectral Adapter (PCSA) to transform the point tokens (serving as graph signal) from the spatial domain to the spectral domain using the spectral basis. Based on the spectral point tokens, our PCSA adopts an effortless shared linear layer to adapt the global and local spectral tokens to meet the downstream tasks. In the spectral domain, adjusting parameters actually modifies the “energy” of each graph frequency for the graph signal, improving the decorrelation of the inner confused tokens when transforming back to the spatial domain. Besides, the information of the generated spectral basis is inherently aware of the unique characteristics of the downstream point clouds, enabling targeted fine-tuning.
@@
Extensive experiments conducted on challenging point cloud datasets across various tasks and data settings demonstrate the effectiveness of our PointGST. For example, when adopting the convincing Point-BERT [85] (resp., Point-MAE [40]) as the baseline, our approach archives notable performance improvement over fully fine-tuning by 3.96%, 1.55%, 2.57%, and 0.7% (resp., 1.72%, 1.90%, 0.11%, and 0.3%) on the ScanObjectNN OBJ_BG, OBJ_ONLY, PB_T50_RS and ModelNet40 datasets, while surprisingly only requires about 2.77% of the total learnable parameters. More importantly, our method even boosts the large-scale pre-trained models PointGPT-L [4], only using 0.6% trainable parameters to achieve 99.48% accuracy on the ScanObjectNN OBJ_BG dataset, outperforming the SOTA method PointGPT-L [4] by 2.28%, establishing the new state-of-the-art. To the best of our knowledge, we are the first to push the performance on this dataset to surpass 99%. We also draw the intuitive trend of performance and trainable parameters in Fig. 1(b).
@@
Our main contributions are summarized as follows: 1) In this paper, we propose a novel Parameter-Efficient Fine-Tuning (PEFT) method for point cloud learning, named PointGST, which innovatively fine-tunes parameters from a fresh perspective, i.e., spectral domain. 2) PointGST introduces the Point Cloud Spectral Adapter (PCSA) for frozen pre-trained backbones, transferring the inner confused tokens from the spatial domain into the spectral domain. It effectively mitigates the confusion among tokens and brings intrinsic information from downstream point clouds to meet the requirement of fine-tuned tasks. 3) Extensive experiments across various datasets demonstrate that our proposed method outperforms all current point cloud-customized PEFT approaches. Remarkably, we achieve state-of-the-art performance on six challenging point cloud datasets while only using about 0.6% trainable parameters, serving as a promising option for fine-tuning the point cloud data.
2 Related works
@@ 2.1 Point Cloud Learning
Constructing structural representations for the point cloud has become a core topic in the computer vision community. To address the irregularity and sparsity of point clouds, the pioneering PointNet [48] adopts shared MLP layers to extract point features independently. The subsequent works further integrate local and global information [49, 56, 52, 37, 61, 72], self-attention mechanisms [96, 19, 78, 77], and various regularization techniques [92, 94, 75], prompting the field.
More recently, self-supervised pre-training methods [85, 50, 81, 36, 98, 16], have gained significant attention in the domain for their remarkable transfer capabilities by learning latent representations from unlabeled data and then fine-tuning on various downstream tasks. Point cloud pre-training approaches can be categorized into contrast-based [81, 1, 6] and reconstruct-based [85, 47, 93, 11, 4, 51] paradigms. In the contrast-based paradigm, PointContrast [81] and CrossPoint [1] extract latent information from different views of the point cloud. Conversely, reconstruction-based methods involve randomly masking point clouds and then using autoencoders to reconstruct the original input. Among them, Point-BERT [85] learns by comparing masked encodings with outputs from a dVAE-based point cloud tokenizer. Point-MAE [47] and Point-M2AE [93] reconstructs original point clouds via an autoencoder. Following this, PointGPT [4] leverages an extractor-generator-based transformer as a GPT-style approach. To tackle data scarcity in 3D representation learning, interest in using cross-modal data for point cloud analysis is growing. For instance, ACT [11] trains 3D representation models using cross-modal pre-trained models as teachers, and ReCon [50, 51] unifies reconstruction and cross-modal contrast modeling further.
Generally, these pre-training approaches typically transfer to the downstream 3D point cloud analysis tasks by fully fine-tuning the pre-trained models. However, as model parameters rapidly increase with scaling up [4, 51], fully fine-tuning incurs high tuning and storage costs and may dilute the pre-training knowledge. This paper seeks to reduce training costs through efficient transfer learning algorithms.
@@ 2.2 Parameter-Efficient Fine-Tuning
Parameter-Efficient Fine-Tuning (PEFT) aims to adopt a trainable module with a few parameters for fine-tuning. It has become a hot research topic in both the NLP [7, 62, 91, 58, 10] and computer vision [26, 35, 86, 67] communities. The mainstream PEFT methods can be roughly categorized into Adapter-based and Prompt-based.
Specifically, the Adapter-based [23, 24, 7, 62, 33] methods usually insert a trainable module between frozen modules to adjust the pre-trained models. Particularly, Adaptformer [7] adds the adapter [23] in the FFN for the visual recognition tasks. The Prompt-based methods [29, 34, 26, 58] usually leverage the additional tokens (i.e., prompts) to the model input. For example, VPT [26], the first prompt-tuning method for vision tasks, prepends additional tunable tokens to the input or hidden layers of ViTs. Apart from paradigms mentioned above, some methods make new attempts, such as directly performing a linear transformation to modulate the features [35] or fine-tuning the bias term in each layer [86].
Recently, some works [88, 64, 99, 15, 32, 63, 60, 14] attempt to explore the PEFT in point cloud tasks. As the earliest method, IDPT [88] applies DGCNN [74] to extract dynamic prompts for different instances instead of the traditional static prompts. Point-PEFT [64] attempts to aggregate local point information during fine-tuning. DA [15] introduces a dynamic aggregation strategy to replace previous static aggregation like mean or max pooling for pre-trained point cloud models. Recently, DAPT [99] proposes an approach combining the dynamic scale adapters with internal prompts as an efficient way of point cloud transfer learning. Although these methods can effectively reduce training costs, they are difficult to achieve the desired performance consistently and generalize poorly. We argue it is attributed to the limitation of fine-tuning in the spatial domain and lack of introducing strong intrinsic information of downstream point cloud. In contrast, we try to fine-tune the parameters in the spectral domain, which effectively mitigates the confusion among tokens and brings intrinsic information about downstream point clouds.

### v2

@@ Abstract
Recently, leveraging pre-training techniques to enhance point cloud models has become a prominent research topic. However, existing approaches typically require full fine-tuning of pre-trained models to achieve satisfactory performance on downstream tasks, which is storage-intensive and computationally demanding. To address this issue, we propose a novel Parameter-Efficient Fine-Tuning (PEFT) method for point cloud, called PointGST (Point cloud Graph Spectral Tuning). PointGST freezes the pre-trained model and introduces a lightweight, trainable Point Cloud Spectral Adapter (PCSA) for fine-tuning parameters in the spectral domain. The core idea is built on two observations: 1) The inner tokens from frozen models might present confusion in the spatial domain; 2) Task-specific intrinsic information is important for transferring the general knowledge to the downstream task. Specifically, PointGST transfers the point tokens from the spatial domain to the spectral domain, effectively de-correlating confusion among tokens by using orthogonal components for separation. Moreover, the generated spectral basis involves intrinsic information about the downstream point clouds, enabling more targeted tuning. As a result, PointGST facilitates the efficient transfer of general knowledge to downstream tasks while significantly reducing training costs. Extensive experiments on challenging point cloud datasets across various tasks demonstrate that PointGST not only outperforms its fully fine-tuning counterpart but also significantly reduces trainable parameters, making it a promising solution for efficient point cloud learning. Moreover, it achieves superior accuracies of 99.48%, 97.76%, and 96.18% on the ScanObjNN OBJ_BG, OBJ_ONLY, and PB_T50_RS datasets, respectively, establishing a new state-of-the-art, while using only 0.67% of the trainable parameters. The code will be released at https://github.com/jerryfeng2003/PointGST.
@@ 1 Introduction
Point cloud learning is a fundamental and highly practical task within the computer vision community, drawing significant attention due to its extensive applications such as autonomous driving [1, 2], 3D reconstruction [3, 4], and embodied intelligence [5], among others. Point clouds are inherently sparse, unordered, and irregular, posing unique challenges to effective analysis.
@@
Nowadays, exploring pre-training techniques to boost the models has become a hot research topic in both natural language processing [6] and computer vision [7], and has been successfully transferred to the point cloud domain [8, 9, 10]. Representative works like Point-BERT [9] and Point-MAE [8] introduce the masked point modeling task for point cloud data, and the subsequent studies [11, 12, 13, 14] propose effective customized pre-training tasks. After pre-training, these methods adopt the classical fully fine-tuning (FFT) strategy, where all models’ parameters are fine-tuned, leading to notable performance improvements and faster convergence compared to training from scratch. However, the FFT strategy incurs significant GPU memory and storage costs due to updating all parameters of the pre-trained models. With the scaling of models [^1] Pre-trained point cloud model parameters increased 30$\times$ in three years, from 22.1M in ECCV 2022 [8] to 657.2M in ECCV 2024 [15]. or increasing needs for fine-tuning enormous new datasets, the demand for storing fine-tuned checkpoints grows, resulting in heightened storage and memory consumption.
@@
The community is well aware of this issue, and to mitigate this challenge, research efforts [16, 17, 18, 19] have been directed towards finding a promising fine-tuning strategy, i.e., Parameter-Efficient Fine-Tuning (PEFT). Their common idea is to freeze the pre-trained models and selectively insert extra learnable parameters into the models. The pioneers are IDPT [16] and DAPT [17], taking an important step toward efficient tuning on point cloud tasks. Specifically, IDPT [16] proposes an instance-aware dynamic prompt tuning to generate a universal prompt for the point cloud data. DAPT [17] generates a dynamic scale for each point token, considering the token significance to the data.
@@
Figure 1: (a) Comparison of existing PEFT methods [17, 16] and our approach. Instead of spatial tuning, we operate in the spectral domain to reduce token confusion and incorporate intrinsic information from downstream point clouds for targeted adaptation. (b) Our PointGST surpasses all prior methods with extremely few trainable parameters, achieving a new record of 99.48% accuracy on ScanObjectNN OBJ_BG and exceeding 99% on this benchmark for the first time.
@@
While the above methods are efficient in training and storage, they still fail to achieve satisfactory performance across various pre-trained models. We attribute this to two factors: First, essentially, these methods are fine-tuned in the spatial domain (Fig. 1(a)) , i.e., the tuning modules process tokens representing localized patches of the point cloud. In the PEFT setting, pre-training tasks are designed to learn general representations of point clouds, without incorporating prior knowledge from downstream tasks. This leads to the inner tokens (features) from frozen pre-trained models struggling to distinguish the fine-grained structures of the point cloud, where such tokens are referred to as inner confused tokens. Specifically, even if two parts of the point cloud have similar geometric structures, their output features from the pre-trained model can differ significantly. Current point cloud PEFT methods address this by merging inner confused tokens with newly introduced learnable modules for downstream task adaptation. However, continuously incorporating confusing features from the frozen parameters during the PEFT task can complicate optimization.
Second, intrinsic structures (e.g., task-specific geometry and relationship of points) of the downstream point clouds are essential for comprehensive analysis [4, 20]. The fixed pre-trained models lack the ability to update their parameters to learn the intrinsic information, relying solely on the representations captured during pre-training. Existing point cloud PEFT methods mainly use the encoded features from the frozen models to generate prompts or feed them into adapters, without explicitly incorporating intrinsic information from downstream point clouds.
@@
To this end, we shift to a fresh perspective: spectral domain fine-tuning. Spectral representations offer significant advantages over spatial domain representations. Specifically, compared to the spatial domain, features in the spectral domain can be easily de-correlated [21, 22], due to the ability of spectral representations to disentangle complex spatial relationships into distinct frequency components. This de-correlation mitigates the confusion among tokens by using orthogonal components (spectral basis) to separate them. Additionally, the spectral basis obtained by the original point cloud offers direct intrinsic information about the downstream tasks to better guide the fine-tuning process. However, effectively and efficiently fine-tuning the parameters in the spectral domain is not a trivial problem, where two crucial concerns need to be considered: 1) what spectral transform to use, and 2) how to inject the intrinsic information into the fixed pre-trained model by using this transform. An ideal transform for point clouds should be easy to construct, preserve intrinsic information, and adapt to the irregular structure of point cloud data. We find that the Graph Fourier Transform (GFT) stands out as particularly suitable. The GFT decomposes confused spatial point tokens into orthogonal graph signals of varying smoothness (the eigenvectors of the Laplace matrix), thus allowing the signal to be separated into orthogonal bases, offering a new perspective for analysis. Moreover, a properly constructed graph provides a natural representation of the irregular point clouds that are adaptive to their structure, and the graph spectral domain can explicitly reveal geometric structures, from basic shapes to fine details, within a compressed space [23].
@@
In this paper, we explore the potential of fine-tuning parameters in the spectral domain. Our approach, illustrated in Fig. 1(a), introduces a novel PEFT method that optimizes a frozen pre-trained point cloud model through Graph Spectral Tuning, which we refer to as PointGST. Specifically, we first construct a series of multi-scale point cloud graphs that enable the calculation of global and local spectral basis, which captures intrinsic information such as the geometry of the original point cloud. Then, we propose a lightweight Point Cloud Spectral Adapter (PCSA) to transform the compressed point tokens (serving as graph signal) from the spatial domain to the spectral domain using the spectral basis. Based on the spectral point tokens, our PCSA adopts an effortless shared linear layer to adapt the global and local spectral tokens to meet the downstream tasks. In the spectral domain, the compressed point tokens are projected onto orthogonal eigenvectors. The goal of spectral domain fine-tuning is to modify the projected coefficients of point tokens within the $n$-dimensional eigenvector space. Due to the bases being orthogonal, the inner confused tokens are de-correlated, thus facilitating the optimization process. Besides, the information of the generated spectral basis is inherently aware of the unique characteristics of the downstream point clouds, enabling targeted fine-tuning.
@@
Extensive experiments conducted on challenging point cloud datasets across various tasks and data settings demonstrate the effectiveness of our PointGST. For example, our method boosts the large-scale pre-trained models PointGPT-L [13], only using 0.67% trainable parameters to achieve 99.48% accuracy on the ScanObjectNN OBJ_BG dataset, establishing the new state-of-the-art. To the best of our knowledge, we are the first to push the performance on this dataset to surpass 99%. We also present the intuitive trend of performance and trainable parameters in Fig. 1(b).
@@
Our main contributions are summarized as follows: 1) We propose a novel Parameter-Efficient Fine-Tuning (PEFT) method for point cloud learning, named PointGST, which innovatively fine-tunes parameters from a fresh perspective, i.e., spectral domain. 2) PointGST introduces the Point Cloud Spectral Adapter (PCSA) for frozen pre-trained backbones, transferring the inner confused tokens from the spatial domain into the spectral domain. It effectively mitigates the confusion among tokens and brings intrinsic information from downstream point clouds to meet the requirements of the fine-tuned tasks. 3) Extensive experiments across various datasets demonstrate that our proposed method outperforms all current point cloud-customized PEFT approaches. Remarkably, we achieve state-of-the-art performance on eight challenging point cloud datasets while only using about 0.67% trainable parameters, serving as a promising option for efficient point cloud fine-tuning.
2 Related work
@@ 2.1 Point Cloud Learning
Constructing structural representations for point clouds is a fundamental problem in computer vision. To address irregularity and sparsity of point clouds, PointNet [24] uses shared MLPs for independent feature extraction, while later works integrate local/global information [25, 26, 27, 28, 29, 30], self-attention [31, 32, 33, 34], and various regularization strategies [35, 36, 10].
Recently, self-supervised pre-training methods [9, 37, 38, 14, 39, 40], have gained significant attention in the domain for their remarkable transfer capabilities by learning latent representations from unlabeled data and then fine-tuning on various downstream tasks. Point cloud pre-training approaches can be categorized into contrast-based [38, 41, 42] and reconstruct-based [9, 8, 12, 11, 13, 15] paradigms. In the contrast-based paradigm, PointContrast [38] and CrossPoint [41] extract latent information from different views of the point cloud. Conversely, reconstruction-based methods involve randomly masking point clouds and then using autoencoders to reconstruct the original input. Among them, Point-BERT [9] learns by comparing masked encodings with outputs from a dVAE-based point cloud tokenizer. Point-MAE [8] and Point-M2AE [12] reconstruct original point clouds via an autoencoder. Following this, PointGPT [13] leverages an extractor-generator-based transformer as a GPT-style approach. To tackle data scarcity in 3D representation learning, interest in using cross-modal data for point cloud analysis is growing. For instance, ACT [11] trains 3D representation models using cross-modal pre-trained models as teachers, and ReCon [37, 15] unifies reconstruction and cross-modal contrast modeling further.
Most pre-trained models are transferred to downstream 3D tasks via full fine-tuning. However, with the growing scale of models [13, 15], full fine-tuning incurs substantial tuning and storage costs and may dilute pre-trained knowledge. This work aims to address these issues through efficient transfer learning algorithms.
@@ 2.2 Parameter-Efficient Fine-Tuning
Parameter-Efficient Fine-Tuning (PEFT) introduces small trainable modules to efficiently adapt large models and has attracted significant attention in both NLP [43, 44, 45, 46, 47] and computer vision [48, 49, 50, 51]. Mainstream PEFT methods are typically Adapter-based [52, 53, 43, 44, 54], which insert adapters between frozen layers or Prompt-based [55, 56, 48, 46], which prepend learnable tokens to model inputs. Other approaches include feature modulation [49] and bias tuning [50].
Recently, some works [16, 18, 17, 19, 57, 58, 59, 60] attempt to explore PEFT in point cloud tasks. As the earliest method, IDPT [16] applies DGCNN [61] to extract dynamic prompts for different instances instead of the traditional static prompts. Point-PEFT [18] attempts to aggregate local point information during fine-tuning. DA [19] introduces a dynamic aggregation strategy to replace previous static aggregation like mean or max pooling for pre-trained point cloud models. DAPT [17] combines the dynamic scale adapters with internal prompts as an efficient way of point cloud transfer learning.
These point cloud methods operate in the spatial domain, where tuning modules handle tokens representing localized patches of the point cloud, each corresponding to a specific point region. We argue that the pre-training tasks prioritize general point cloud representations without prior downstream tasks, leading to frozen pre-trained models struggling to distinguish fine-grained point cloud structures in the spatial domain.